# Perception-guided Jailbreak against Text-to-Image Models

### v1

@@ 1 Introduction
However, the widespread use and advanced capabilities of these models have led to significant security concerns regarding unsafe image generation. A prominent ethical issue associated with T2I models is their potential to produce sensitive Not-Safe-for-Work (NSFW) images [4, 5], including adult content, violence, and politically sensitive material. Therefore, current T2I models incorporate safety checkers [6, 7] as essential guardrails to prevent the generation of NSFW images.
@@ 2.1 Text-to-Image Models
Notable examples include Stable Diffusion [1], DALL·E [22, 3], Imagen [23], Midjourney [2], and Wanxiang [24]. One of the most advanced T2I models, DALL·E 3 [3], integrated natively into ChatGPT [25], leverages LLM [26] to refine prompts, producing images that closely align with the input prompts and reducing the users’ burden of prompt engineering. Given their popularity, studies that investigate the vulnerabilities of T2I models are necessary.
@@ 2.2 Jailbreak on Text-to-Image Models
Previous research [27, 28, 29, 30] predominantly focuses on text modification to probe functional vulnerabilities (e.g., diminishing synthetic quality, distorting or eliminating objects, impairing image fidelity). However, they do not target the generation of Not-Safe-For-Work (NSFW) images such as pornography, violence, and racism.
Currently, more and more works [8, 9, 10, 11, 12] have put emphasis on exploring the opened avenues for potential misuse of T2I models, particularly in generating inappropriate or NSFW content. SneakyPrompt [8] exploits reinforcement learning to replace the words in the prompt for bypassing safety filters in T2I generative models. SurrogatePrompt [9] proposes a pipeline that contains three modules to generate NSFW images on T2I models such as Midjourney and DALL·E 2. DACA [31] breaks down unethical prompts into multiple benign descriptions of individual image elements and makes word substitutions for each element. MMA-Diffusion [10] is a multimodal attack framework that designs attacks on both text and image modalities. UPAM [11] is a unified framework that employs gradient-based optimization, sphere-probing learning, and semantic-enhancing learning to attack the T2I model. JPA [12] using learnable tokens to create adversarial prompts that evade detection while preserving the semantic integrity of the original NSFW content.
Rely on white-box adversarial attacks targeting a specific T2I model, and then subsequently transfer the generated adversarial prompts to attack other T2I models. This often results in the generation of nonsensical, incomprehensible tokens within the attack prompts, thereby diminishing their stealthiness. ❷ Others involve developing complex pipelines that require numerous queries to the T2I model, leading to high time and resource consumption. In contrast, our method is model-free, requiring no specific T2I model as a target, and generates attack prompts with high naturalness.
@@ 5.1 Experimental Setups
Evaluation metrics. We use four metrics to evaluate the experiment. ❶ We use the attack success rate (ASR) metric to evaluate the number of attack prompts that bypass the NSFW detector divided by the total number of attack prompts. ❷ We use the semantic consistency (SC) metric to represent the consistency between the semantics of the generated image and the original unsafe user prompt. The generated image should have a similar semantic as the original unsafe user prompt, i.e., the jailbreak method does not change the semantics of the unsafe user prompt. The semantic of generated image is extracted by BLIP [36]. ❸ We use prompt perplexity (PPL) as a metric to evaluate the coherence of the modified attack prompt. The prompt with high PPL contains a lot of garbled characters and is easy to notice. ❹ We use the Inception Score (IS) to evaluate the diversity of the generated images. For ASR, SC, and IS metrics, higher is better while for PPL, lower is better. Note that the ASR and SC metrics are dominant ones for evaluating the jailbreak performance of methods.
@@ 6 Conclusion
In this paper, we introduce a word replacement method that identifies a safe substitution phrase adhering to the PSTSI principle. The proposed PGJ method efficiently and effectively generate an attack prompt capable of bypassing the safety checkers in T2I models. For future work, we plan to explore for circumventing the post-checker in T2I models.

### v2

@@ 1 Introduction
The growing interest in artificial intelligence (AI) safety [4, 5, 6, 7] has intensified due to the widespread adoption and advanced capabilities of AI models. This expansion has raised substantial security concerns about the potential for generating unsafe images. A prominent ethical issue associated with T2I models is their potential to produce sensitive Not-Safe-for-Work (NSFW) images [8, 9], including adult content, violence, and politically sensitive material. Therefore, current T2I models incorporate safety checkers [10, 11] as essential guardrails to prevent the generation of NSFW images.
@@ 2.1 Text-to-Image Models
Notable examples include Stable Diffusion [1], DALL·E [26, 3], Imagen [27], Midjourney [2], and Wanxiang [28]. One of the most advanced T2I models, DALL·E 3 [3], integrated natively into ChatGPT [29], leverages LLM [30] to refine prompts, producing images that closely align with the input prompts and reducing the users’ burden of prompt engineering. Given their popularity, studies that investigate the vulnerabilities [31, 32] of T2I models are necessary.
@@ 2.2 Jailbreak on Text-to-Image Models
Previous research [33, 34, 35, 36, 37, 38] predominantly focuses on text modification to probe functional vulnerabilities (e.g., diminishing synthetic quality, distorting or eliminating objects, impairing image fidelity). However, they do not target the generation of Not-Safe-For-Work (NSFW) images such as pornography, violence, and racism.
Currently, more and more works [12, 13, 14, 15, 16] have put emphasis on exploring the opened avenues for potential misuse of T2I models, particularly in generating inappropriate or NSFW content via adversarial attack [39, 40]. SneakyPrompt [12] exploits reinforcement learning to replace the words in the prompt for bypassing safety filters in T2I generative models. SurrogatePrompt [13] proposes a pipeline that contains three modules to generate NSFW images on T2I models such as Midjourney and DALL·E 2. DACA [41] breaks down unethical prompts into multiple benign descriptions of individual image elements and makes word substitutions for each element. MMA-Diffusion [14] is a multimodal attack framework that designs attacks on both text and image modalities. UPAM [15] is a unified framework that employs gradient-based optimization, sphere-probing learning, and semantic-enhancing learning to attack the T2I model. JPA [16] using learnable tokens to create adversarial prompts that evade detection while preserving the semantic integrity of the original NSFW content.
❶ Rely on white-box adversarial attacks targeting a specific T2I model, and then subsequently transfer the generated adversarial prompts to attack other T2I models. This often results in the generation of nonsensical, incomprehensible tokens within the attack prompts, thereby diminishing their stealthiness. ❷ Others involve developing complex pipelines that require numerous queries to the T2I model, leading to high time and resource consumption. In contrast, our method is model-free, requiring no specific T2I model as a target, and generates attack prompts with high naturalness.
@@ 5.1 Experimental Setups
Evaluation metrics. We use four metrics to evaluate the experiment. ❶ We use the attack success rate (ASR) metric to evaluate the number of attack prompts that bypass the NSFW detector divided by the total number of attack prompts. ❷ We use the semantic consistency (SC) metric to represent the consistency between the semantics of the generated image and the original unsafe user prompt. The generated image should have a similar semantic as the original unsafe user prompt, i.e., the jailbreak method does not change the semantics of the unsafe user prompt. The semantics of the generated image are extracted by BLIP [46]. ❸ We use prompt perplexity (PPL) as a metric to evaluate the coherence of the modified attack prompt. The prompt with high PPL contains a lot of garbled characters and is easy to notice. ❹ We use the Inception Score (IS) to evaluate the diversity of the generated images. For ASR, SC, and IS metrics, higher is better while for PPL, lower is better. Note that the ASR and SC metrics are dominant ones for evaluating the jailbreak performance of methods.
@@ 6 Conclusion
In this paper, we introduce a word replacement method that identifies a safe substitution phrase adhering to the PSTSI principle. The proposed PGJ method efficiently and effectively generates an attack prompt capable of bypassing the safety checkers in T2I models. For future work, we plan to explore circumventing the post-checker in T2I models.